Circular No. 6948

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SGR 1627–41 P. Woods, M. Kippen, and Jan van Paradijs, University of Alabama

in Huntsville; C. Kouveliotou and M. McCollough, Universities Space Research Association; and K. Hurley, Space Sciences Laboratory, communicate: “BATSE Earth-limb considerations limit the possible position of this object to declinations along the IPN annulus reported by Hurley et al. (below), which are between −43^{o} and −49^{o}. This is slightly outside the BATSE 1σ error circle reported on IAUC 6944. However, due to the soft spectrum of the source, systematic errors were larger than usual. A preliminary search through the MOST supernova remnant (SNR) catalogue (Whiteoak and Green 1996, A.Ap. Suppl. 118, 329) reveals only one candidate within the error box reported here, G337.0–0.1. The peculiar nonthermal-structure core reported for G337.0–0.1 is entirely within the IPN annulus width. We strongly encourage follow-up observations of the SNR source.”

Hurley, on behalf of the Ulysses Gamma-Ray Burst Team; and Kouveliotou, Kippen, and Woods, on behalf of the Gamma-Ray Observatory BATSE team, report: “We have obtained an IPN annulus for this soft γray repeater (IAUC 6944). Its center is at α = 22^{h}01^{m}40^{s}, δ = −10^{o}09′.8 (equinox 2000.0), and its radius is 76^{o}.769 ± 0^{o}.028. Considerable further refinement of this position is possible. An image may be found at

http://ssl.berkeley.edu/ipn3/sgr1627-41/.”

COMET C/1998 K2 (LINEAR)

Ephemeris extension to IAUC 6915 (elements MPC 31893):

1998 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
June 26 15^{h}39.^{m}63 −42^{o}51.^{′}5 1.561 2.443 141.^{o}9 14.^{o}9 13.3
July 6 15 17.49 −45 36.6 1.630 2.411 129.8 18.9 13.4
16 14 59.58 −47 48.4 1.721 2.383 118.8 21.9 13.5
26 14 46.79 −49 39.1 1.827 2.361 108.9 24.0 13.5
Aug. 5 14 39.28 −51 20.2 1.941 2.343 100.2 25.2 13.6
15 14 36.74 −53 00.1 2.057 2.331  92.4 25.7 13.7
25 14 38.81 −54 43.7 2.171 2.324  85.7 25.7 13.8
Sept. 4 14 45.19 −56 34.1 2.281 2.323  79.7 25.3 14.0

Visual m_{1} estimates: May 31.27 UT, 13.1 (C. S. Morris, Lockwood Valley, CA, 0.5-m reflector); June 2.64, 12.6 (M. Mattiazzo, Wallaroo, South Australia, 0.20-m reflector); 17.54, 12.7 (Mattiazzo).

1998 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green